# Reduced Order Modelling, Dynamical Systems, Deep Learning, Uncertainty Quantification journal: Journal name

### v2

@@ Abstract
Reduced order models (ROMs) play a critical role in fluid mechanics by providing low-cost predictions, making them an attractive tool for engineering applications. However, for ROMs to be widely applicable, they must not only generalise well across different regimes, but also provide a measure of confidence in their predictions. While recent data-driven approaches have begun to address nonlinear reduction techniques to improve predictions in transient environments, challenges remain in terms of robustness and parametrisation. In this work, we present a nonlinear reduction strategy specifically designed for transient flows that incorporates parametrisation and uncertainty quantification. Our reduction strategy features a variational auto-encoder (VAE) that uses variational inference for confidence measurement. We use a latent space transformer that incorporates recent advances in attention mechanisms to predict dynamical systems. Attention’s versatility in learning sequences and capturing their dependence on external parameters enhances generalisation across a wide range of dynamics. Prediction, coupled with confidence, enables more informed decision making and addresses the need for more robust models. In addition, this confidence is used to cost-effectively sample the parameter space, improving model performance a priori across the entire parameter space without requiring evaluation data for the entire domain.
@@ 1 Introduction
The integration of machine learning with physics has fostered a mutually beneficial relationship, driving significant advancements in computational modelling. This synergy has given rise to the development of both physics-informed and physics-inspired models. Physics-informed models, like PINNs [1], explicitly incorporate physical laws into their structure, ensuring that solutions adhere to known scientific principles [2]. On the other hand, physics-inspired models, such as Hamiltonian and Lagrangian networks, are based on the conceptual framework of classical mechanics and other physical theories. These models draw inspiration from the rich mathematical structures found in physics, using concepts like energy conservation, action minimization, and symmetries to inform their architecture and training procedures. By leveraging these fundamental principles, physics-inspired models are able to encode essential physical behaviours into their design, often leading to enhanced generalization and more physically plausible predictions. For example, Hamiltonian networks use the principles of Hamiltonian mechanics to model the dynamics of a system, while Lagrangian networks leverage the idea of least action to optimize their predictions  [3, 4]. These approaches allow for the modelling of complex physical systems by inherently incorporating dynamical properties and constraints, without requiring explicit equations for every interaction, thus capturing the essence of physical systems in a computationally efficient manner.
@@
One notable application of these advanced modelling techniques is Reduced Order Modelling (ROM). ROMs aim to offer computationally efficient alternatives to large, data-intensive models, which is especially valuable for real-time simulations and control in critical systems [10]. By simplifying the system’s representation, ROMs improve interpretability and transparency, key factors in applications where understanding the underlying mechanics is just as crucial as making accurate predictions. Traditional ROM approaches, such as Proper Orthogonal Decomposition (POD)[11], achieve dimensionality reduction by projecting data onto a lower-dimensional manifold based on the overall contribution to the energy norm. An alternative to POD, Dynamic Mode Decomposition (DMD)[12], focuses on decomposing complex systems into the most dominant time-evolving modes. While both POD and DMD, along with their variations [13, 14], have proven effective in identifying coherent structures within data, they are limited by their linear nature, which restricts their ability to fully capture nonlinear dynamics. Consequently, there has been a growing shift toward nonlinear methods [15, 16]. In this context, machine learning techniques has been increasingly used to both identify the reduced basis [17] and to represent the dynamics within the lower-dimensional manifold [18]. A recent advancement in this area is the machine-learning closure model for POD, referred to as CD-ROM [19], where the projection error orthogonal to the POD basis is addressed by a historical term predicted through a Neural ODE. This methodology offers a more accurate representation of the system’s dynamics, significantly enhancing the performance of reduced order models in capturing the complexities of nonlinear systems.
@@
Koopman theory provides a robust mathematical framework for dealing with complex dynamical systems by assuming the existence of an infinite-dimensional space in which the dynamics are linear. This concept has been applied in various machine learning implementations [20], but it requires expanding the number of dimensions towards infinity, which is often impractical and different from the Reduced Order Model (ROM) framework that our work seeks to integrate. Work combining ROMs with Koopman theory has shown impressive results on complex dynamical problems such as the Navier-Stokes equations [21, 22].
@@
However, these methods have two major limitations. First, they need to be parametrised, or in other words learn and generalise across different operating conditions, referred to in this paper as ’parameters’. Second, they are inherently deterministic, which limits their ability to account for uncertainty in predictions. To address these limitations, we emphasise the importance of adopting a probabilistic framework. Such a framework allows for the quantification of uncertainty, which is critical for providing confidence levels in predictions. A probabilistic approach also offers improved robustness to model inaccuracies and support for risk-aware decision making. These benefits are critical for fostering confidence and enabling the integration of machine learning models into model-based design.
@@ 2 Components of UP-dROM
where $F$ is a nonlinear operator governing the spatial and temporal evolution of the system, $\phi$ represents the state variables and $\xi$ the parameters governing the dynamical behaviour of the system. The objective is to develop a reduced-order model that accurately represents the dynamics of the given partial differential equation. Considering the high dimensionality of the problem and the inherent presence of coherent structures (modes) in most dynamical systems, we employ dimensionality reduction techniques to both preserve the coherence and reduce the system’s dimensionality. To effectively analyse and capture these space-time coherent structures, we propose a space-time strategy that compresses spatial information while simultaneously models the temporal dependencies within this lower-dimensional manifold.
@@ 2.1 Dimensionality reduction
In this work, we reduce the spatial component of the system using a Variational Autoencoder (VAE), which facilitates the compression of high-dimensional physical data into a lower-dimensional nonlinear manifold. The choice of a VAE over a traditional autoencoder is motivated by its ability to incorporate uncertainty quantification, which is a critical aspect of our approach, as discussed in Section 4. The VAE is particularly advantageous because it introduces a probabilistic framework, enabling the model to not only capture the underlying structure of the data but also quantify the uncertainty in the learned representations [24]. This probabilistic nature allows for more reliable inference, especially in situations where uncertainty plays a significant role in the prediction or system behaviour. Variational Autoencoders offer the advantage of learning continuous, disentangled latent representations. This structured latent space improves the robustness of inference and enhances generalization beyond the training distribution, especially when compared to traditional Auto-Encoders. The ability to generalize more effectively is a key factor driving the widespread use of VAEs in generative AI frameworks [31].
Moreover, VAEs and Autoencoders (AEs) have been shown to outperform traditional methods in compressing data from complex physical systems [21, 24, 32, 33, 25]. This is because VAEs amd AEs do not impose strict orthogonality constraints, which is a key limitation in traditional linear compression techniques, such as Proper Orthogonal Decomposition (POD), PCA [11, 34]. In contrast, VAEs and AEs offer the flexibility to relax this constraint, making them better suited for capturing intricate nonlinear structures within the data. Furthermore, the absence of a linearity constraint enables VAEs to generalize more effectively to unseen data distributions. This ability to generalize is particularly important when dealing with real-world physical systems, where the underlying dynamics may vary across different regimes or boundary conditions.
@@ 2.2 Temporal evolution
To capture the time component of these structures within the latent space, several strategies can be employed. These include utilizing a Koopman operator [22], solving the governing equations analytically using Galerkin projection in conjunction with a deep-learning closure model [19], or applying Long Short-Term Memory (LSTM) networks to capture both short-term and long-term temporal dependencies [35]. In this study, we propose the use of Transformers, similar to those employed in large language models (LLMs) [30]. This choice is driven by two primary considerations: (i) the attention mechanism in Transformers serves as an effective memory component, facilitating the model’s ability to retain relevant information over time, and (ii) the combination of self-attention and cross-attention mechanisms enables the model to capture complex and nonlinear dependencies, including interactions with external variables. This approach enhances the model’s generalization capabilities and overall robustness. The resulting architecture is illustrated in Figure 1.
@@ 2.3 Model training
Transformer and auto-encoder are trained together to promote a symbiotic relationship between the two elements, thus promoting an organic modal representation. The total loss is calculated considering a prediction horizon $h$ and a lookback window $q$. The appendix provides further details on the selection of hyperparameters across different test cases, which were chosen through grid search to balance computational efficiency and model accuracy. The loss function is defined as:
@@
where $\mathcal{E}$ is the encoding process, $\mathcal{D}$ the decoding process, KLD the Kullback-Leibler divergence and $\hat{z}$ the latent variables. In addition, $\lambda$ and $\beta$ are regularisation terms. The first term of this loss promotes a mirroring process between encoder and decoder, similar to a classical projection. The Kullback-Leibler divergence term is related to the use of the VAE and is explained in section 4.
@@ 2.4 Validation of UP-dROM – application to flow around a cylinder
The model is validated by predicting the evolution of a flow around a blunt object, specifically a cylinder. The flow dynamics are governed by the incompressible Navier-Stokes equations :
@@
Where, $u$ is the velocity field, $p$ is the pressure, and $Re$ is the Reynolds number. The data is generated using a solver based on the Immersed Boundary Method [36]. This solver utilizes two grids: a regular Eulerian grid that spans the entire flow domain and a Lagrangian grid that conforms to the boundary of the bluff body, enabling accurate representation of the obstacle-fluid interaction. Once generated, only the Eulerian grid data of the flow field is used; the model is not explicitly informed of the bluff body’s position or shape and must learn it implicitly from the data.
The data set provided to the model includes all Eulerian grid points and starts from the stationary unstable solution and captures the transient dynamics leading to the stable limit cycle solution. It includes velocities $u$ and $v$ in $x$ and $y$ directions respectively, stacked in a state vector $\Phi$. The data set is divided into a training and a test data set. In this test case, including two crucial phases, to ensure equal representation of transients and the limit cycle in both the training and test data sets, we split the data in half by assigning even time steps to the training set and odd time steps to the test set.
The flow behaviour is primarily influenced by the Reynolds number $R_{e}$, which is treated here as the external variable or parameter $\xi$. The Reynolds number is a dimensionless quantity that characterizes the ratio of inertial forces to viscous forces in a fluid flow and is defined as:
@@
where, $\rho$ is the fluid density, $U$ is the characteristic velocity, $L$ is the characteristic length (such as the length of the bluff body), and $\mu$ is the dynamic viscosity of the fluid. For this particular bluff body, when $R_{e}$ remains below its critical value, the flow closely resembles a stationary solution, with minimal deviations over time. However, beyond this critical Reynolds number, vortex shedding and unsteady flow dynamics emerge in the wake of the cylinder, leading to a more complex and time-dependent behaviour. Figure 2 shows the base flow provided to the model at the initial time $t_{0}$ and a representative frame at a later time $t$ when the limit cycle is established. The model should be able to capture the transition dynamics and the stable limit cycle dynamics.
@@
The transition from the base flow to the stable limit cycle is captured through the evolution of the kinetic energy signal. In Figure 2, we compare the inferred kinetic energy to the ground truth, followed by a forecasting window that demonstrates the model’s performance. The kinetic energy can be computed from the state vector $\phi$, which includes both the velocity fields $u$ and $v$ over time $t$, with a spatial discretization given by $n_{xy}$. The formula for kinetic energy is given by:
@@
This metric effectively averages the spatial components in a way that is physically meaningful. By eliminating the spatial dimensions, it enables us to concentrate on time-dependent signals, making it a valuable visualization tool for this study. The results show that the model performs comparably to similar architectures and strategies that use delay embedding to represent the system dynamics in the latent space [22, 19].
A key decision in our approach is the preference for a Transformer model over other time-embedding architectures, such as Long Short-Term Memory (LSTM) networks. This choice is informed by recent advancements in embedding memory mechanisms, which significantly enhance the model’s predictive capabilities [29, 37, 25, 38]. Specifically, Transformers leverage the attention mechanism, which allows the model to capture long-range dependencies in the time series data more effectively than traditional methods. In addition to self-attention, the incorporation of cross-attention mechanisms in Transformers enables the model to respond adaptively to external variables or influences, providing a more comprehensive understanding of the system dynamics. This architecture choice also supports the broader objective of improving model generalization. By enabling the model to learn relationships and interactions between the data and an external variable such as $R_{e}$, Transformers are particularly well-suited to address the challenges of generalizing across different conditions, thereby enhancing the robustness and versatility of the model in diverse physical system scenarios.
@@
Consider the example of flow around a bluff body, as shown in Figure 2, which serves as a simple dynamical system with a varying Reynolds number, as depicted in Figure 3. As the Reynolds number changes, the system undergoes a bifurcation and transitions into an unstable state. This is evidenced by the steady-state branch bifurcating into oscillatory branches, with unsteady von Kármán vortex street patterns emerging in the flow domain. For the case of a flow 2D around a bluff body considered here, we considered an elliptical obstacle, with an aspect-ratio $\beta=1.35$. This aspect-ratio sets the bifurcation at a critical Reynolds Number $R_{ec}\approx 60$.
@@
To test whether the performance of UP-dROM extends to a larger dimensional parameter space, the aspect ratio of the bluff body was added as a second parameter. The aspect-ratio, as defined in Figure 6, characterizes the shape of the bluff body and significantly influences the flow behaviour. An aspect-ratio smaller than 1 indicates that the obstacle is taller than wide, creating a stronger disruption in the flow. This increased blockage leads to earlier vortex shedding and a bifurcation occurring at a lower Reynolds number. In contrast, an aspect-ratio greater than 1 corresponds to an elongated body, similar to an airfoil, which disturbs the flow less significantly. In this case, the onset of vortex shedding and bifurcation is delayed, occurring at a higher Reynolds number due to the more streamlined geometry. This effect is illustrated in Figure 6, where different aspect ratios result in distinct flow regimes, even though the Reynolds number remains unchanged.
Figure 6: Aspect ratio definition (a) and $u$ velocity field at $R_{e}=50$ for aspet-ratio $\beta=1.5$ (b), and $\beta=0.75$ (c).
Performance with a two-variable parametrisation $(R_{e},\beta)$ can be visualised using a pointwise scaled MSE, as shown in Figure 7, which ensures that each grid point is equally weighted in the evaluation. The scaled MSE, computed on test data, is obtained locally using the magnitude of local variations. The resulting heatmap provides a a-posteriori perspective on the model’s ability to generalise and its overall performance. The signals shown on the map represent the evolution of the kinetic energy for that configuration over time. The model performs well on the training sets, demonstrating its ability to learn and predict. In particular, its interpolation results are strong. However, similar behaviour to the single parameter case is observed. UP-dROM is able to distinguish between stable and limit-cycle dynamics in the two-parameter regime. However, the accuracy of the prediction deteriorates when the model is used in extrapolation mode.

### v3

@@ Abstract
Reduced order models (ROMs) play a critical role in fluid mechanics by providing low-cost predictions, making them an attractive tool for engineering applications. However, for ROMs to be widely applicable, they must not only generalise well across different regimes, but also provide a measure of confidence in their predictions. While recent data-driven approaches have begun to address nonlinear reduction techniques to improve predictions in transient environments, challenges remain in terms of robustness and parametrisation.
In this work, we present a nonlinear reduction strategy specifically designed for transient flows that incorporates parametrisation and uncertainty quantification. Our reduction strategy features a variational autoencoder(VAE) that uses variational inference for confidence measurement. We use a latent space transformer that incorporates recent advances in attention mechanisms to predict dynamical systems. Attention’s versatility in learning sequences and capturing their dependence on external parameters enhances generalisation across a wide range of dynamics. Prediction, coupled with confidence, enables more informed decision-making and addresses the need for more robust models.
In addition, this confidence is used to cost-effectively sample the parameter space, improving model performance a priori across the entire parameter space without requiring evaluation data for the entire domain.
@@ 1 Introduction
The integration of machine learning with physics has fostered a mutually beneficial relationship, driving significant advancements in computational modelling. This synergy has given rise to the development of both physics-informed and physics-inspired models. Physics-informed models, like PINNs [1], explicitly incorporate physical laws into their structure, ensuring that solutions adhere to known scientific principles [2]. On the other hand, physics-inspired models, such as Hamiltonian and Lagrangian networks, are based on the conceptual framework of classical mechanics and other physical theories. These models draw inspiration from the rich mathematical structures found in physics, using concepts like energy conservation, action minimisation, and symmetries to inform their architecture and training procedures. By leveraging these fundamental principles, physics-inspired models are able to encode essential physical behaviours into their design, often leading to enhanced generalisation and more physically plausible predictions. For example, Hamiltonian networks use the principles of Hamiltonian mechanics to model the dynamics of a system, while Lagrangian networks leverage the idea of least action to optimise their predictions  [3, 4]. These approaches allow for the modelling of complex physical systems by inherently incorporating dynamical properties and constraints, without requiring explicit equations for every interaction, thus capturing the essence of physical systems in a computationally efficient manner.
@@
One notable application of these advanced modelling techniques is Reduced Order Modelling (ROM). ROMs aim to offer computationally efficient alternatives to large, data-intensive models, which is especially valuable for real-time simulations and control in critical systems [10]. By simplifying the system representation, ROMs improve interpretability and transparency, key factors in applications where understanding the underlying mechanics is just as crucial as making accurate predictions. Traditional ROM approaches, such as Proper Orthogonal Decomposition (POD)[11], achieve dimensionality reduction by projecting data onto a lower-dimensional manifold based on the overall contribution to the energy norm. An alternative to POD, Dynamic Mode Decomposition (DMD)[12], focuses on decomposing complex systems into the most dominant time-evolving modes. While both POD and DMD, along with their variations [13, 14], have proven effective in identifying coherent structures within data, they are limited by their linear nature, which restricts their ability to fully capture nonlinear dynamics. Consequently, there has been a growing shift towards nonlinear methods [15, 16]. In this context, machine learning techniques have been increasingly used to both identify the reduced basis [17] and to represent the dynamics within the lower-dimensional manifold [18]. A recent advancement in this area is the machine-learning closure model for POD, referred to as CD-ROM [19], where the projection error orthogonal to the POD basis is addressed by a historical term predicted through a Neural ODE. This methodology offers a more accurate representation of the system’s dynamics, significantly enhancing the performance of reduced order models in capturing the complexities of nonlinear systems.
@@
Koopman theory provides a robust mathematical framework for dealing with complex dynamical systems by assuming the existence of an infinite-dimensional space in which the dynamics are linear. This concept has been applied in various machine learning implementations [20], but it requires increasing the dimensionality towards infinity, which is often impractical and different from the Reduced Order Model (ROM) framework that our work seeks to integrate. Work combining ROMs with Koopman theory has shown impressive results on complex dynamical problems such as the Navier-Stokes equations [21, 22].
@@
However, these methods have two major limitations. First, they need to be parametrised, or in other words learn and generalise across different operating conditions, referred to in this paper as ’parameters’. Second, they are inherently deterministic, which limits their ability to account for uncertainty in predictions. To address these limitations, we emphasise the importance of adopting a probabilistic framework. Such a framework allows for the quantification of uncertainty, which is critical for providing confidence levels in predictions. A probabilistic approach also offers improved robustness to model inaccuracies and support for risk-aware decision-making. These benefits are critical for fostering confidence and enabling the integration of machine learning models into model-based design.
@@ 2 Components of UP-dROM
where $F$ is a nonlinear operator governing the spatial and temporal evolution of the system, $\phi$ represents the state variables and $\xi$ the parameters governing the dynamical behaviour of the system. The objective is to develop a reduced-order model that accurately represents the dynamics of the given partial differential equation. Considering the high dimensionality of the problem and the inherent presence of coherent structures (modes) in most dynamical systems, we employ dimensionality reduction techniques to both preserve the coherence and reduce the system’s dimensionality. To effectively analyse and capture these space-time coherent structures, we propose a space-time strategy that compresses spatial information while simultaneously modelling the temporal dependencies within this lower-dimensional manifold.
@@ 2.1 Dimensionality reduction
In this work, we reduce the spatial component of the system using a Variational Autoencoder (VAE), which facilitates the compression of high-dimensional physical data into a lower-dimensional nonlinear manifold. The choice of a VAE over a traditional autoencoder is motivated by its ability to incorporate uncertainty quantification, which is a critical aspect of our approach, as discussed in Section 4. The VAE is particularly advantageous because it introduces a probabilistic framework, enabling the model to not only capture the underlying structure of the data but also quantify the uncertainty in the learned representations [24]. This probabilistic nature allows for more reliable inference, especially in situations where uncertainty plays a significant role in prediction or system behaviour. Variational Autoencoders offer the advantage of learning continuous, disentangled latent representations. This structured latent space improves the robustness of inference and enhances generalisation beyond the training distribution, especially when compared to traditional Auto-Encoders. The ability to generalise more effectively is a key factor driving the widespread use of VAEs in generative AI frameworks [31].
Moreover, VAEs and Autoencoders (AEs) have been shown to outperform traditional methods in compressing data from complex physical systems [21, 24, 32, 33, 25]. This is because VAEs and AEs do not impose strict orthogonality constraints, which is a key limitation in traditional linear compression techniques, such as Proper Orthogonal Decomposition (POD), PCA [11, 34]. In contrast, VAEs and AEs offer the flexibility to relax this constraint, making them better suited for capturing intricate nonlinear structures within the data. Furthermore, the absence of a linearity constraint enables VAEs to generalise more effectively to unseen data distributions. This ability to generalise is particularly important when dealing with real-world physical systems, where the underlying dynamics may vary across different regimes or boundary conditions.
@@ 2.2 Temporal evolution
To capture the time component of these structures within the latent space, several strategies can be employed. These include utilizing a Koopman operator [22], solving the governing equations analytically using Galerkin projection in conjunction with a deep-learning closure model [19], or applying Long Short-Term Memory (LSTM) networks to capture both short-term and long-term temporal dependencies [35]. In this study, we propose the use of Transformers, similar to those employed in large language models (LLMs) [30]. This choice is driven by two primary considerations: (i) the attention mechanism in Transformers serves as an effective memory component, facilitating the model’s ability to retain relevant information over time, and (ii) the combination of self-attention and cross-attention mechanisms enables the model to capture complex and nonlinear dependencies, including interactions with external variables. This approach enhances the model’s generalisation capabilities and overall robustness. The resulting architecture is illustrated in Figure 1.
@@ 2.3 Model training
Transformer and autoencoderare trained together to promote a symbiotic relationship between the two elements, thus fostering an organic modal representation. The total loss is calculated considering a prediction horizon $h$ and a lookback window $q$. The appendix provides further details on the selection of hyperparameters across different test cases, which were chosen through grid search to balance computational efficiency and model accuracy. The loss function is defined as:
@@
where $\mathcal{E}$ is the encoding process, $\mathcal{D}$ the decoding process, KLD the Kullback-Leibler divergence and $\hat{z}$ the latent variables. In addition, $\lambda$ and $\beta$ are regularisation coefficients. The first term of this loss promotes a mirroring process between encoder and decoder, similar to a classical projection. The Kullback-Leibler divergence term is related to the use of the VAE and is explained in Section 4.
@@ 2.4 Validation of UP-dROM – application to flow around a cylinder
The model is validated by predicting the evolution of flow around a blunt object, specifically a cylinder. The flow dynamics are governed by the incompressible Navier-Stokes equations :
@@
Where $u$ is the velocity field, $p$ is the pressure, and $Re$ is the Reynolds number. The data are generated using a solver based on the Immersed Boundary Method [36]. This solver utilises two grids: a regular Eulerian grid that spans the entire flow domain and a Lagrangian grid that conforms to the boundary of the bluff body, enabling accurate representation of the obstacle-fluid interaction. Once generated, only the Eulerian grid data of the flow field are used; the model is not explicitly informed of the bluff body’s position or shape and must learn it implicitly from the data.
The dataset provided to the model includes all Eulerian grid points and starts from the stationary unstable solution and captures the transient dynamics leading to the stable limit cycle solution. It includes velocities $u$ and $v$ in the $x$ and $y$ directions, respectively, stacked in a state vector $\Phi$. The dataset is divided into a training and a test sets. In this test case, including two crucial phases, to ensure equal representation of the transients and the limit cycle in both the training and test sets, we split the data in half by assigning even time steps to the training set and odd time steps to the test set.
The flow behaviour is primarily influenced by the Reynolds number $R_{e}$, which is treated here as the external variable or parameter $\xi$. The Reynolds number is a dimensionless quantity that characterises the ratio of inertial forces to viscous forces in a fluid flow and is defined as:
@@
where $\rho$ is the fluid density, $U$ is the characteristic velocity, $L$ is the characteristic length (such as the length of the bluff body), and $\mu$ is the dynamic viscosity of the fluid. For this particular bluff body, when $R_{e}$ remains below its critical value, the flow closely resembles a stationary solution, with minimal deviations over time. However, beyond this critical Reynolds number, vortex shedding and unsteady flow dynamics emerge in the wake of the cylinder, leading to a more complex and time-dependent behaviour. Figure 2 shows the base flow provided to the model at the initial time $t_{0}$ and a representative frame at a later time $t$ when the limit cycle is established. The model should be able to capture the transition dynamics and the stable limit cycle dynamics.
@@
The transition from the base flow to the stable limit cycle is captured through the evolution of the kinetic energy signal. In Figure 2, we compare the inferred kinetic energy to the ground truth, followed by a forecasting window that demonstrates the model’s performance. The kinetic energy can be computed from the state vector $\phi$, which includes both the velocity fields $u$ and $v$ over time $t$, with a spatial discretisation given by $n_{xy}$. The formula for kinetic energy is given by:
@@
This metric effectively averages the spatial components in a way that is physically meaningful. By eliminating the spatial dimensions, it enables us to concentrate on time-dependent signals, making it a valuable visualisation tool for this study. The results show that the model performs comparably to similar architectures and strategies that use delay embedding to represent the system dynamics in the latent space [22, 19].
A key decision in our approach is the preference for a Transformer model over other time-embedding architectures, such as Long Short-Term Memory (LSTM) networks. This choice is informed by recent advancements in embedding memory mechanisms, which significantly enhance the model’s predictive capabilities [29, 37, 25, 38]. Specifically, Transformers leverage the attention mechanism, which allows the model to capture long-range dependencies in the time-series data more effectively than traditional methods. In addition to self-attention, the incorporation of cross-attention mechanisms in Transformers enables the model to respond adaptively to external variables or influences, providing a more comprehensive understanding of the system dynamics. This architecture choice also supports the broader objective of improving model generalisation. By enabling the model to learn relationships and interactions between the data and an external variable such as $R_{e}$, Transformers are particularly well-suited to address the challenges of generalizing across different conditions, thereby enhancing the robustness and versatility of the model in diverse physical system scenarios.
@@
Consider the example of flow around a bluff body, as shown in Figure 2, which serves as a simple dynamical system with a varying Reynolds number, as depicted in Figure 3. As the Reynolds number changes, the system undergoes a bifurcation and transitions into an unstable state. This is evidenced by the steady-state branch bifurcating into oscillatory branches, with unsteady von Kármán vortex street patterns emerging in the flow domain. For the case of 2D flow around a bluff body considered here, we considered an elliptical obstacle with an aspect-ratio $\beta=1.35$. This aspect-ratio sets the bifurcation at a critical Reynolds Number $R_{ec}\approx 60$.
@@
To test whether the performance of UP-dROM extends to a larger dimensional parameter space, the aspect ratio of the bluff body was added as a second parameter. The aspect-ratio, as defined in Figure 6, characterises the shape of the bluff body and significantly influences the flow behaviour. An aspect-ratio smaller than 1 indicates that the obstacle is taller than wide, creating a stronger disruption in the flow. This increased blockage leads to earlier vortex shedding and a bifurcation occurring at a lower Reynolds number. In contrast, an aspect-ratio greater than 1 corresponds to an elongated body, similar to an airfoil, which disturbs the flow less significantly. In this case, the onset of vortex shedding and bifurcation is delayed, occurring at a higher Reynolds number due to the more streamlined geometry. This effect is illustrated in Figure 6, where different aspect ratios result in distinct flow regimes, even though the Reynolds number remains unchanged.
Figure 6: Aspect ratio definition (a) and $u$ velocity field at $R_{e}=50$ for aspect-ratio $\beta=1.5$ (b), and $\beta=0.75$ (c).
Performance with a two-variable parametrisation $(R_{e},\beta)$ can be visualised using a pointwise scaled MSE, as shown in Figure 7, which ensures that each grid point is equally weighted in the evaluation. The scaled MSE, computed on test data, is obtained locally using the magnitude of local variations. The resulting heatmap provides a a-posteriori perspective on the model’s ability to generalise and its overall performance. The signals shown on the heatmap represent the evolution of the kinetic energy for that configuration over time. The model performs well on the training sets, demonstrating its ability to learn and predict. In particular, its interpolation results are strong. However, similar behaviour to the single parameter case is observed. UP-dROM is able to distinguish between stable and limit-cycle dynamics in the two-parameter regime. However, the accuracy of the prediction deteriorates when the model is used in extrapolation mode.